andy and brian together have $40, brian and chris have $62 and andy and chris have $58. how much money do the three of them have together ?
a $160 b $80 c $40 d $100 e $120

Answer by Theo(13342)

a = money that andy has.
b = money that brian has.
c = money that chris has.

a + b = 40
b + c = 62
a + c = 58

use your first equation to solve for b in terms of a.
you get b = 40-a

use your third equation to solve for c in terms of a.
you get c = 58-a

substitute for b and c in your second equation to get:

b + c = 62 becomes:
40-a + 58-a = 62
combine like terms to get:
98 - 2a = 62
add 2a to both sides of this equation and subtract 62 from both sides of this equation to get:
98 - 62 = 2a
combine like terms to get:
36 = 2a
divide both sides of this equation by 2 to get:
18 = a

substitute for a in the first equation to get:

a + b = 40 becomes:
18 + b = 40
solve for b to get:
b = 40-18 = 22

substitute for a in the third equation to get:

a + c = 58 becomes:
18 + c = 58
solve for c to get:
c = 58-18 = 40

your solution should be:
a = 18
b = 22
c = 40

substitute in your original equations to see if these values are good.

original equations are:

a + b = 40
b + c = 62
a + c = 58

These become:

18 + 22 = 40 (true)
22 + 40 = 62 (true)
18 + 40 = 58 (true)

Values are good.

Answer is:

The money that all 3 of them have together is:

18 + 22 + 40 = $80.00

That would be selection b.